# Prototypical Contrastive Learning-based CLIP Fine-tuning for Object Re-identification

### v2

@@ Abstract
This work aims to adapt large-scale pre-trained vision-language models, such as contrastive language-image pre-training (CLIP), to enhance the performance of object re-identification (Re-ID) across various supervision settings. Although prompt learning has enabled a recent work named CLIP-ReID to achieve promising performance, the underlying mechanisms and the necessity of prompt learning remain unclear due to the absence of semantic labels in Re-ID tasks. In this work, we first analyze the role prompt learning in CLIP-ReID and identify its limitations. Based on our investigations, we propose a simple yet effective approach to adapt CLIP for supervised object Re-ID. Our approach directly fine-tunes the image encoder of CLIP using a prototypical contrastive learning (PCL) loss, eliminating the need for prompt learning. Experimental results on both person and vehicle Re-ID datasets demonstrate the competitiveness of our method compared to CLIP-ReID. Furthermore, we extend our PCL-based CLIP fine-tuning approach to unsupervised scenarios, where we achieve state-of-the-art performance.
@@ 1 Introduction
Recently, contrastive language-image pre-training (CLIP) [clip] and other pre-trained vision-language models [Jia2021, Du2022] have attracted great attention in vision community. CLIP is trained on a dataset of 400 million text-image pairs collected from the internet. Through unsupervised cross-modal contrastive learning from the large-scale dataset, it is capable of learning diverse visual and language semantic concepts and acquiring remarkable transfer abilities. As a result, CLIP has been successfully adapted to various downstream vision tasks, including zero-shot/few-shot recognition [coop, cocoop], object detection [Zhong2022RegionCLIP, Wu2023CORA], semantic segmentation [Luo2023SegCLIP], and more.
@@
Figure 1: t-SNE [tsne] visualization of randomly selected 7 IDs from MSMT17. (a) shows that the text centroids learned in CLIP-ReID stage-1 are pretty close to the image centroids, which reveals their implicit equivalence. (b) shows that PCL is also able to learn high-quality feature space only with image centroids. Best view with color.
@@
Various techniques, such as learning additional adaptation layers [gao2021clip] and learning textual or visual prompts [coop, cocoop, Jia2022, Khattak2023], have been developed to facilitate the adaptation of CLIP to specific downstream tasks. Among these adaptation techniques, prompt learning has gained popularity due to its superior performance and low computational cost. Both textual [coop, cocoop] and multi-modal [Khattak2023] prompt learning use class names from ground-truth labels to form input text descriptions for recognition or classification tasks. The inclusion of fixed class names helps the CLIP model transfer textual semantics to visual concepts, resulting in great robustness to noise [Wu2023] and remarkable generalization ability to unseen data [coop].
@@
Unfortunately, in person/vehicle re-identification, class names do not exist as the class labels represent ID indexes and lack semantic meaning. This poses a challenge when employing prompt learning techniques to adapt pre-trained vision-language models. A recent work called CLIP-ReID [clipreid] addresses this issue by introducing a two-stage strategy. In the first stage, it learns a set of textual prompts for each ID while keeping the CLIP model fixed. Then, in the second stage, the learned prompts are utilized to fine-tune the image encoder of CLIP using a proposed image-to-text cross-entropy loss, in conjunction with the commonly used ID loss and triplet loss [luo2019trick]. This approach has demonstrated impressive performance in supervised Re-ID, seeming like highlighting the potential of prompt learning in adapting CLIP to the Re-ID task.
@@
However, upon investigating the mechanisms that enable effective textual prompt learning in CLIP-ReID, we have made the following speculations: 1) Unlike CoOP [coop] that utilizes fixed class names to transfer textual semantics, the prompt learning in CLIP-ReID essentially learns a textual feature centroid for each ID, as shown in Figure 1. 2) The image-to-text cross-entropy loss introduced in CLIP-ReID serves as a centroid-based loss [Wen2016, Attias2017proxy], attracting images of the same ID towards their respective textual centroids. Inspired by these findings, we propose to utilize a prototypical contrastive learning (PCL) loss [Li2021] to directly fine-tune the CLIP’s image encoder without the need of prompt learning. Our PCL loss leverages the ID centroids of up-to-date visual features for fine-tuning, preventing from the disturbance caused by outdated textual centroids and potential misalignment between textual and visual features. Experimental results demonstrate that simply fine-tuning CLIP with a single PCL loss performs competitively with CLIP-ReID [clipreid], achieving significantly higher performance compared to fine-tuning CLIP with ID loss and triplet loss.
@@
The aforementioned fine-tuning of CLIP is performed under full supervision, where ID labels are available. In this work, we also aim to adapt CLIP to unsupervised object Re-ID. The dominant approach for unsupervised Re-ID is clustering-based, which generates pseudo labels through clustering and then use the labels to learn a Re-ID model iteratively. The iterative learning scheme and the varying number of pseudo labels at different iterations make it challenging to learn prompts similar to CLIP-ReID [clipreid].
@@
Fortunately, our PCL loss-based CLIP fine-tuning fits well within this framework. Considering that state-of-the-art (SoTA) methods such as ClusterContrast [clustercontrast], CAP [cap], and O2CAP [O2CAP] already adopt PCL-wise losses, we can simply replace their backbone with the image encoder of CLIP and fine-tune the models using their own PCL-wise losses. However, directly replacing and fine-tuning the CLIP image encoder leads to a divergence issue due to the instability of training vision transformers under unsupervised setting. To address this issue, we adopt the patch projection layer frozen trick [Chen2021ViT-ssl], which allows our model to converge effectively and achieve significantly higher performance compared to SoTA methods.
@@ 2.1 Object Re-identification
In the past decade, fully supervised object Re-ID has witnessed remarkable advancements, primarily driven by the utilization of deep convolutional neural networks (CNNs) like ResNet-50 [he2016deep], in combination with the use of ID loss and triplet loss [hermans2017defense, luo2019trick]. Numerous CNN-based methods have been developed, leveraging multi-granularity features [sun2018beyond], human semantics [zhang2019dsa], attention mechanisms [chen2019abd], etc. More recently, transformer architectures [transreid, AAformer, LA-Transformer] and pre-trained vision-language models [clip, clipreid] have also been employed to boost Re-ID performance.
Purely unsupervised Re-ID has also witnessed significant progress in recent years. A majority of research is clustering-based, which involves generating pseudo labels through clustering and iteratively learning a Re-ID model using these labels. Previous methods have made efforts to refine noisy pseudo labels [RLCC, PPLR, O2CAP], leverage contrastive learning losses [cap, O2CAP, clustercontrast], and design network architectures [transreidssl, TMGF]. To the best of our knowledge, the utilization of pre-trained vision-language models has not been explored in unsupervised Re-ID.
@@ 2.2 CLIP and Prompt Learning
Contrastive language-image pre-training (CLIP) [clip] is a vision-language model that undergoes pre-training on millions of text-image pairs available on the Internet. It jointly trains a text encoder and an image encoder using two directional InfoNCE losses, which are extensively used in contrastive learning [Oord2019CPC, Chen2020SimCLR, he2020momentum]. Beneficial from the knowledge learned from the large-scale dataset, CLIP has demonstrated remarkable generalization ability and has been effectively transferred to various downstream vision tasks, including zero-shot recognition [coop, cocoop], object detection [Zhong2022RegionCLIP, Wu2023CORA], semantic segmentation [Luo2023SegCLIP], among others.
Prompt learning [coop, cocoop, Jia2022, Khattak2023] has gained popularity in effectively transferring CLIP to specific downstream tasks. As a pioneering work, CoOp [coop] introduces the concept of learning text prompts, which significantly outperformed hand-crafted prompts in CLIP adaptation. Building upon this, CoCoOp [cocoop] further enhances generalizability by allowing text prompts to be conditioned on inputs, VPT [Jia2022] and MaPLe [Khattak2023] extend the learning to visual and multi-modal prompts. In both textual [coop, cocoop] and multi-modal [Khattak2023] prompt learning, the class names of ground-truth labels are utilized to form input text descriptions, contributing remarkable generalization ability to unseen data [coop] and enhancing robustness to noise [Wu2023].
However, in object re-identification, class names do not exist as the labels are represented by ID indexes. In order to adapt CLIP to Re-ID tasks, CLIP-ReID [clipreid] introduces a two-stage strategy that first learns a set of textual prompts for each ID and then employs a prompt-based loss to fine-tune CLIP, resulting in impressive performance. Nevertheless, when using our prototypical contrastive learning loss to fine-tune the CLIP image encoder, this prompt learning strategy brings almost no improvement for both fully supervised and unsupervised Re-ID tasks. These results suggest that prompt learning may not be necessary for achieving strong performance in Re-ID tasks and our PCL loss-based fine-tuning approach can be a promising alternative.
@@ 2.3 Prototypical Contrastive Learning
Prototypical contrastive learning (PCL) [Li2021] is a type of contrastive learning method that operates at the cluster level. In this approach, each class or cluster is represented by a prototype, which is a central representation of the instances within that cluster. The primary objective of PCL is to attract each instance towards its own cluster’s prototype while simultaneously pushing it away from the prototypes of other clusters. By leveraging the cluster-level information, prototypical contrastive learning can effectively capture local semantic structures, resulting in more effective learning compared to conventional instance-level contrastive learning methods commonly used in uni-modal and multi-modal unsupervised representation learning tasks[Chen2020SimCLR, he2020momentum, clip]. In recent years, prototypical contrastive learning has been successfully applied to unsupervised Re-ID tasks. Various PCL-wise losses, such as leveraging camera-aware prototypes [cap] or online associated prototypes [O2CAP], have been developed, leading to significant improvements in performance. In this work, we explore the use of the PCL loss for fine-tuning CLIP in both fully supervised and unsupervised Re-ID tasks.
@@ 3 Revisit CLIP-ReID
We first revisit the CLIP-ReID method [clipreid]. It proposes a two-stage strategy to adapt CLIP [clip] to supervised object Re-ID. Let us denote the pre-trained text and image encoders of CLIP as $\mathcal{T}(\cdot)$ and $\mathcal{I}(\cdot)$, respectively. These two encoders map a sequence of text prompt tokens and an image into features $\mathbf{f}^{T}$ and $\mathbf{f}^{I}$ in a joint embedding space.
In the first stage, CLIP-ReID learns a set of ID-specific text tokens while freezing both the text and image encoders like CoOP [coop]. For an image $i$ with an ID label $y_{i}=c$ ($c\in\{1,\cdots,C\}$), the paired text sequence input into $\mathcal{T}(\cdot)$ is designed as “A photo of $X]^{c}_{1}[X]^{c}_{2}\cdots[X]^{c}_{M}$ person/vehicle”. Here, $X]^{c}_{m}$ ($m\in\{1,\cdots,M\}$) is a learnable token with the same dimension as word embedding, $C$ denotes the total number of IDs and $M$ is the nubmer of learnable tokens. The text tokens are learned by optimizing the sum of the following image-to-text and text-to-image losses:
@@
In the second stage, CLIP-ReID fine-tunes the image encoder of CLIP while fixing both the learned prompts and the text encoder. The image encoder is optimized with respect to a ID loss $\mathcal{L}_{id}$ and a triplet loss $\mathcal{L}_{tri}$ that are widely used in supervised Re-ID [luo2019trick], together with an image-to-text cross-entropy loss defined as below:
@@
Discussion. In the CLIP-ReID framework, each ID class $c$ is associated with a textual feature $\mathbf{f}_{c}^{T}$, which is obtained by feeding a sequence containing the learned ID-specific prompts into the text encoder. The aforementioned loss $\mathcal{L}{i2tce}$ serves to attract the visual feature of an image towards its corresponding textual feature while pushing it away from textual features of other classes. This implies that each class’s textual feature is well-aligned with its visual feature centroid. However, achieving such alignment is not guaranteed due to the following reasons. Firstly, although the training of CLIP encourages the alignment of textual and visual features in a joint embedding space, there may still be slight separation due to the modality gap [Liang2022, Shi2023]. Secondly, even if the two encoders are initially well-aligned in CLIP, fine-tuning the image encoder while keeping the text encoder fixed may introduce misalignment issues. Therefore, while the loss $\mathcal{L}{i2tce}$ significantly improves performance compared to using only the ID loss and triplet loss for fine-tuning CLIP, the presence of misalignment issues can have a negative impact on the fine-tuning process, resulting in sub-optimal performance.
@@ 4 PCL-based CLIP Fine-tuning
Based on our understanding of the mechanisms behind the effectiveness and potential limitations of prompt learning in CLIP-ReID [clipreid], we propose a direct fine-tuning approach for CLIP using prototypical contrastive learning (PCL). In this work, we first apply PCL-based CLIP fine-tuning for the supervised Re-ID task and then extend our approach to unsupervised settings. By leveraging the benefits of PCL, we aim to improve the performance of CLIP in adapting to Re-ID tasks while bypassing the need for prompt learning.
@@ 4.1 Fine-tuning for Supervised Re-ID
Prototypical contrastive learning. Prototypical contrastive learning has been widely used in unsupervised representation learning [Li2021] and unsupervised Re-ID [cap, O2CAP, clustercontrast] tasks. The objective of PCL is to bring an instance closer to its cluster centroid while pushing it away from the centroids of other clusters. By this means, PCL learns to distinguish between different clusters and capture the underlying structure of the data in an unsupervised manner.
@@
Discussion. In contrast to CLIP-ReID [clipreid] that utilizes a loss to attract or repel an image towards its textual feature centroid, our PCL loss operates directly on the visual features, eliminating the need for alignment between textual and visual features. Experimental results show that fine-tuning CLIP with a single PCL loss achieves competitive performance compared to CLIP-ReID, which employs three losses for fine-tuning. Furthermore, while prototypical contrastive learning has also been successfully applied to supervised image classification [Khosla2020], the ID loss and triplet loss remain dominant in supervised Re-ID methods. Our study indicates that the PCL loss is more effective than these two losses for adapting CLIP to supervised Re-ID tasks.

### v3

@@ Abstract
This work aims to adapt large-scale pre-trained vision-language models, such as contrastive language-image pre-training (CLIP), to enhance the performance of object re-identification (Re-ID) across various supervision settings. Although prompt learning has enabled a recent work named CLIP-ReID to achieve promising performance, the underlying mechanisms and the necessity of prompt learning remain unclear due to the absence of semantic labels in Re-ID tasks. In this work, we first analyze the role prompt learning in CLIP-ReID and identify its limitations. Based on our investigations, we propose a simple yet effective approach to adapt CLIP for supervised object Re-ID. Our approach directly fine-tunes the image encoder of CLIP using a prototypical contrastive learning (PCL) loss, eliminating the need for prompt learning. Experimental results on both person and vehicle Re-ID datasets demonstrate the competitiveness of our method compared to CLIP-ReID. Furthermore, we extend our PCL-based CLIP fine-tuning approach to unsupervised scenarios, where we achieve state-of-the-art performance. Code is available at https://github.com/RikoLi/PCL-CLIP.
@@ 1 Introduction
Recently, contrastive language-image pre-training (CLIP) [34] and other pre-trained vision-language models [17, 8] have attracted great attention in vision community. CLIP is trained on a dataset of 400 million text-image pairs collected from the internet. Through unsupervised cross-modal contrastive learning from the large-scale dataset, it is capable of learning diverse visual and language semantic concepts and acquiring remarkable transfer abilities. As a result, CLIP has been successfully adapted to various downstream vision tasks, including zero-shot/few-shot recognition [54, 53], object detection [51, 46], semantic segmentation [31], and more.
@@
Figure 1: t-SNE [40] visualization of randomly selected 7 IDs from MSMT17. (a) shows that the text centroids learned in CLIP-ReID stage-1 are pretty close to the image centroids, which reveals their implicit equivalence. (b) shows that PCL is also able to learn high-quality feature space only with image centroids. Best view with color.
@@
Various techniques, such as learning additional adaptation layers [10] and learning textual or visual prompts [54, 53, 18, 20], have been developed to facilitate the adaptation of CLIP to specific downstream tasks. Among these adaptation techniques, prompt learning has gained popularity due to its superior performance and low computational cost. Both textual [54, 53] and multi-modal [20] prompt learning use class names from ground-truth labels to form input text descriptions for recognition or classification tasks. The inclusion of fixed class names helps the CLIP model transfer textual semantics to visual concepts, resulting in great robustness to noise [45] and remarkable generalization ability to unseen data [54].
@@
Unfortunately, in person/vehicle re-identification, class names do not exist as the class labels represent ID indexes and lack semantic meaning. This poses a challenge when employing prompt learning techniques to adapt pre-trained vision-language models. A recent work called CLIP-ReID [25] addresses this issue by introducing a two-stage strategy. In the first stage, it learns a set of textual prompts for each ID while keeping the CLIP model fixed. Then, in the second stage, the learned prompts are utilized to fine-tune the image encoder of CLIP using a proposed image-to-text cross-entropy loss, in conjunction with the commonly used ID loss and triplet loss [29]. This approach has demonstrated impressive performance in supervised Re-ID, seeming like highlighting the potential of prompt learning in adapting CLIP to the Re-ID task.
@@
However, upon investigating the mechanisms that enable effective textual prompt learning in CLIP-ReID, we have made the following speculations: 1) Unlike CoOP [54] that utilizes fixed class names to transfer textual semantics, the prompt learning in CLIP-ReID essentially learns a textual feature centroid for each ID, as shown in Figure 1. 2) The image-to-text cross-entropy loss introduced in CLIP-ReID serves as a centroid-based loss [44, 32], attracting images of the same ID towards their respective textual centroids. Inspired by these findings, we propose to utilize a prototypical contrastive learning (PCL) loss [24] to directly fine-tune the CLIP’s image encoder without the need of prompt learning. Our PCL loss leverages the ID centroids of up-to-date visual features for fine-tuning, preventing from the disturbance caused by outdated textual centroids and potential misalignment between textual and visual features. Experimental results demonstrate that simply fine-tuning CLIP with a single PCL loss performs competitively with CLIP-ReID [25], achieving significantly higher performance compared to fine-tuning CLIP with ID loss and triplet loss.
@@
The aforementioned fine-tuning of CLIP is performed under full supervision, where ID labels are available. In this work, we also aim to adapt CLIP to unsupervised object Re-ID. The dominant approach for unsupervised Re-ID is clustering-based, which generates pseudo labels through clustering and then use the labels to learn a Re-ID model iteratively. The iterative learning scheme and the varying number of pseudo labels at different iterations make it challenging to learn prompts similar to CLIP-ReID [25].
@@
Fortunately, our PCL loss-based CLIP fine-tuning fits well within this framework. Considering that state-of-the-art (SoTA) methods such as ClusterContrast [7], CAP [41], and O2CAP [42] already adopt PCL-wise losses, we can simply replace their backbone with the image encoder of CLIP and fine-tune the models using their own PCL-wise losses. However, directly replacing and fine-tuning the CLIP image encoder leads to a divergence issue due to the instability of training vision transformers under unsupervised setting. To address this issue, we adopt the patch projection layer frozen trick [5], which allows our model to converge effectively and achieve significantly higher performance compared to SoTA methods.
@@ 2.1 Object Re-identification
In the past decade, fully supervised object Re-ID has witnessed remarkable advancements, primarily driven by the utilization of deep convolutional neural networks (CNNs) like ResNet-50 [14], in combination with the use of ID loss and triplet loss [16, 29]. Numerous CNN-based methods have been developed, leveraging multi-granularity features [38], human semantics [48], attention mechanisms [3], etc. More recently, transformer architectures [15, 56, 36] and pre-trained vision-language models [34, 25] have also been employed to boost Re-ID performance.
Purely unsupervised Re-ID has also witnessed significant progress in recent years. A majority of research is clustering-based, which involves generating pseudo labels through clustering and iteratively learning a Re-ID model using these labels. Previous methods have made efforts to refine noisy pseudo labels [47, 6, 42], leverage contrastive learning losses [41, 42, 7], and design network architectures [30, 23]. To the best of our knowledge, the utilization of pre-trained vision-language models has not been explored in unsupervised Re-ID.
@@ 2.2 CLIP and Prompt Learning
Contrastive language-image pre-training (CLIP) [34] is a vision-language model that undergoes pre-training on millions of text-image pairs available on the Internet. It jointly trains a text encoder and an image encoder using two directional InfoNCE losses, which are extensively used in contrastive learning [39, 4, 13]. Beneficial from the knowledge learned from the large-scale dataset, CLIP has demonstrated remarkable generalization ability and has been effectively transferred to various downstream vision tasks, including zero-shot recognition [54, 53], object detection [51, 46], semantic segmentation [31], among others.
Prompt learning [54, 53, 18, 20] has gained popularity in effectively transferring CLIP to specific downstream tasks. As a pioneering work, CoOp [54] introduces the concept of learning text prompts, which significantly outperformed hand-crafted prompts in CLIP adaptation. Building upon this, CoCoOp [53] further enhances generalizability by allowing text prompts to be conditioned on inputs, VPT [18] and MaPLe [20] extend the learning to visual and multi-modal prompts. In both textual [54, 53] and multi-modal [20] prompt learning, the class names of ground-truth labels are utilized to form input text descriptions, contributing remarkable generalization ability to unseen data [54] and enhancing robustness to noise [45].
However, in object re-identification, class names do not exist as the labels are represented by ID indexes. In order to adapt CLIP to Re-ID tasks, CLIP-ReID [25] introduces a two-stage strategy that first learns a set of textual prompts for each ID and then employs a prompt-based loss to fine-tune CLIP, resulting in impressive performance. Nevertheless, when using our prototypical contrastive learning loss to fine-tune the CLIP image encoder, this prompt learning strategy brings almost no improvement for both fully supervised and unsupervised Re-ID tasks. These results suggest that prompt learning may not be necessary for achieving strong performance in Re-ID tasks and our PCL loss-based fine-tuning approach can be a promising alternative.
@@ 2.3 Prototypical Contrastive Learning
Prototypical contrastive learning (PCL) [24] is a type of contrastive learning method that operates at the cluster level. In this approach, each class or cluster is represented by a prototype, which is a central representation of the instances within that cluster. The primary objective of PCL is to attract each instance towards its own cluster’s prototype while simultaneously pushing it away from the prototypes of other clusters. By leveraging the cluster-level information, prototypical contrastive learning can effectively capture local semantic structures, resulting in more effective learning compared to conventional instance-level contrastive learning methods commonly used in uni-modal and multi-modal unsupervised representation learning tasks[4, 13, 34]. In recent years, prototypical contrastive learning has been successfully applied to unsupervised Re-ID tasks. Various PCL-wise losses, such as leveraging camera-aware prototypes [41] or online associated prototypes [42], have been developed, leading to significant improvements in performance. In this work, we explore the use of the PCL loss for fine-tuning CLIP in both fully supervised and unsupervised Re-ID tasks.
@@ 3 Revisit CLIP-ReID
We first revisit the CLIP-ReID method [25]. It proposes a two-stage strategy to adapt CLIP [34] to supervised object Re-ID. Let us denote the pre-trained text and image encoders of CLIP as $\mathcal{T}(\cdot)$ and $\mathcal{I}(\cdot)$, respectively. These two encoders map a sequence of text prompt tokens and an image into features $\mathbf{f}^{T}$ and $\mathbf{f}^{I}$ in a joint embedding space.
In the first stage, CLIP-ReID learns a set of ID-specific text tokens while freezing both the text and image encoders like CoOP [54]. For an image $i$ with an ID label $y_{i}=c$ ($c\in\{1,\cdots,C\}$), the paired text sequence input into $\mathcal{T}(\cdot)$ is designed as “A photo of $X]^{c}_{1}[X]^{c}_{2}\cdots[X]^{c}_{M}$ person/vehicle”. Here, $X]^{c}_{m}$ ($m\in\{1,\cdots,M\}$) is a learnable token with the same dimension as word embedding, $C$ denotes the total number of IDs and $M$ is the nubmer of learnable tokens. The text tokens are learned by optimizing the sum of the following image-to-text and text-to-image losses:
@@
In the second stage, CLIP-ReID fine-tunes the image encoder of CLIP while fixing both the learned prompts and the text encoder. The image encoder is optimized with respect to a ID loss $\mathcal{L}_{id}$ and a triplet loss $\mathcal{L}_{tri}$ that are widely used in supervised Re-ID [29], together with an image-to-text cross-entropy loss defined as below:
@@
Discussion. In the CLIP-ReID framework, each ID class $c$ is associated with a textual feature $\mathbf{f}_{c}^{T}$, which is obtained by feeding a sequence containing the learned ID-specific prompts into the text encoder. The aforementioned loss $\mathcal{L}{i2tce}$ serves to attract the visual feature of an image towards its corresponding textual feature while pushing it away from textual features of other classes. This implies that each class’s textual feature is well-aligned with its visual feature centroid. However, achieving such alignment is not guaranteed due to the following reasons. Firstly, although the training of CLIP encourages the alignment of textual and visual features in a joint embedding space, there may still be slight separation due to the modality gap [26, 37]. Secondly, even if the two encoders are initially well-aligned in CLIP, fine-tuning the image encoder while keeping the text encoder fixed may introduce misalignment issues. Therefore, while the loss $\mathcal{L}{i2tce}$ significantly improves performance compared to using only the ID loss and triplet loss for fine-tuning CLIP, the presence of misalignment issues can have a negative impact on the fine-tuning process, resulting in sub-optimal performance.
@@ 4 PCL-based CLIP Fine-tuning
Based on our understanding of the mechanisms behind the effectiveness and potential limitations of prompt learning in CLIP-ReID [25], we propose a direct fine-tuning approach for CLIP using prototypical contrastive learning (PCL). In this work, we first apply PCL-based CLIP fine-tuning for the supervised Re-ID task and then extend our approach to unsupervised settings. By leveraging the benefits of PCL, we aim to improve the performance of CLIP in adapting to Re-ID tasks while bypassing the need for prompt learning.
@@ 4.1 Fine-tuning for Supervised Re-ID
Prototypical contrastive learning. Prototypical contrastive learning has been widely used in unsupervised representation learning [24] and unsupervised Re-ID [41, 42, 7] tasks. The objective of PCL is to bring an instance closer to its cluster centroid while pushing it away from the centroids of other clusters. By this means, PCL learns to distinguish between different clusters and capture the underlying structure of the data in an unsupervised manner.
@@
Discussion. In contrast to CLIP-ReID [25] that utilizes a loss to attract or repel an image towards its textual feature centroid, our PCL loss operates directly on the visual features, eliminating the need for alignment between textual and visual features. Experimental results show that fine-tuning CLIP with a single PCL loss achieves competitive performance compared to CLIP-ReID, which employs three losses for fine-tuning. Furthermore, while prototypical contrastive learning has also been successfully applied to supervised image classification [21], the ID loss and triplet loss remain dominant in supervised Re-ID methods. Our study indicates that the PCL loss is more effective than these two losses for adapting CLIP to supervised Re-ID tasks.